OpenStudy (anonymous):

Let f(x) = –4x + 7 and g(x) = 10x – 6, find f(g(x)).

OpenStudy (anonymous):

i got to -40x^2 + 94x - 42 and now im stuck

OpenStudy (anonymous):

not quite right

OpenStudy (anonymous):

\[f(g(x))\] does not mean f times g. it means f OF g so compute as follows \[f(g(x))=f(10x-6)=-4(10x-6)+7\] then algebra

OpenStudy (anonymous):

when its (f o g)(x), u plug in the value of g(x) for x in f(x)